OpenStudy (anonymous):

how to graph inequalities: how can you solve inequalities by the method dotted line and solid line positive or negative and shading the region.... im confused

OpenStudy (anonymous):

medals will be be awarded..

OpenStudy (anonymous):

also please dont use any tables

OpenStudy (anonymous):

\[< and > \] should be dotted lines. \[\le and \ge\] Should be solid lines. Tables are one of the easiest ways to do it.

OpenStudy (anonymous):

thank you so much but do you know if the slope is postivie or negative should i shade above or under to find that region

OpenStudy (anonymous):

Well if the equation is say y<-mx+b the slope is neg if it's just y<mx+b than the slope is pos.

OpenStudy (anonymous):

wait i mean about the shading thing when you have 2 equations

OpenStudy (anonymous):

and the position of the line when the slope is either positive or negative

OpenStudy (anonymous):

Give me an example

OpenStudy (anonymous):

such as y<3x+5

OpenStudy (anonymous):



OpenStudy (anonymous):

help..

OpenStudy (anonymous):

Do you know how to graph a normal equation?

OpenStudy (anonymous):

this needs to be done by hand and simple its solved with 2 inequalities

OpenStudy (anonymous):

thanks though and yes

OpenStudy (anonymous):

Well to determine which side is shaded you plug in any point and if it's true then that side is to be shaded if it's not true then shade the other side.

OpenStudy (anonymous):

its just supposed to be regional where you shade not exact... thats why i asked when its positive or negative is the line supposed to be slanted to the right or left

OpenStudy (anonymous):

pos to the right neg to the left

OpenStudy (anonymous):

ok thanks so much for your help

OpenStudy (anonymous):

No problem if you'd like to see a 'hand written' one I could graph one by hand virtually.

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

make sure this has 2 equations

OpenStudy (anonymous):

i mean inequalities

OpenStudy (anonymous):

Well give me an example of what you're looking for.

OpenStudy (anonymous):

\[y \ge x+and 3y <-6x+6\]

OpenStudy (anonymous):

thanks @jonjenkins7653

OpenStudy (anonymous):

no prob

OpenStudy (anonymous):

im sorry the first inquality is y is greater than or equal to x+2

OpenStudy (anonymous):

\[y \ge x+2\] is the same as graphing y=x+2 The start at 2 and go up 1 over (to the right) 1 \[3y < -6x+6\] You must first get the y by itself by dividing by 3 which you end up getting \[y < -2x + 2\] So to graph it start at 2 and go Down 2 over (to the right)1 Just a sec while I graph just the lines.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

So here is just the lines graphed I used virtual graph paper to graph these lines. Blue is \[y \ge x+2\] Orange is \[y<-2x+2\]

OpenStudy (anonymous):

all right so far so good but arent you going to shade above or below the line to see what regions they cross at

OpenStudy (anonymous):

Now for the shading. Take the first inequality \[y \ge x+2\] And plug in (0,0) for the x and y this helps you determine which side to shade. \[0 \ge 0+2\] \[0 \ge 2\] That is not true so don't shade the half the has the point (0,0)

OpenStudy (anonymous):

so youre saying if something is not true you dont shade?

OpenStudy (anonymous):

Then you don't shade that half.

OpenStudy (anonymous):

ok thanks so much for your time and effort thanks for helping

OpenStudy (anonymous):

;)

OpenStudy (anonymous):

Here is the final graph.